investment of 4000 in 3 months with interest of 6.1 what was the earned amount?
Answer by checkley77(12844)

P=(1+r/n)^t/n
Where P=principle, r=rate (in decimal form),n=number of periods per year. t=number of years.
4,000(1+.061/4)^1/4
4,000(1+.01525)^.25
4,000(1.01525)^.25
4,000*1.00379=$4,015.16 ans.